1. simplify
(sqrt -64)+ 6i^2,555

2. Simplify.
(2-i^22)^3

Answer by Alan3354(69443)

1. simplify
(sqrt -64)+ 6i^2,555
sqrt(-64) = 8i
i^2555 = i^2552*i^3 = 1*-i = -i
--> 8i - 6i
= 2i
================
2. Simplify.
(2-i^22)^3
i^22 = i^20*i^2 = 1*-1 = -1
--> (2+1)^3
= 27